Personal Record

Shortened Version

Complete Version: http://www.hedorfer.it/docs/acurr/acurrcom-ENG.pdf

Markus M. Hedorfer

Urban and Regional Planning — Geospatial Information Systems Via Ca’ Rossa 93, I–30174 Venezia–Mestre VE Telephone: ++39–041–2668833, Telefax: ++39–041–2668834 Electronic mail: mm@hedorfer.it WWW–home–page: http://www.hedorfer.it

11th October 2004

Personal Data

Birth Place:  Karlsruhe, FR Germany
Birth Date:  2nd January 1965
Residence:  Venezia–Mestre, Italy
Marital Status: married
Citizenship:  German
Tax-Code:  Italy–HDR MKS 65A02 Z112N
VAT-Code:  IT02966250272

Scholar and Professional Carreer

Chronology

1984 Attainment of the German general matriculation standard at the Municipal Secondary School (Städtisches Gymnasium) of Hechingen, FR Germany. Matriculation at the Venice University of Architecture (Istituto Universitario di Architettura di Venezia, IUAV) at the Graduation Course in Urban and Regional Planning (Pianificazione Territoriale e Urbanistica) for the academic year 1984–1985. 1986 Six–week’s practical course at the Regionalverband (Region) of Neckar– Alb at Tübingen, FR Germany. 1994 Completion of the programme of studies with its focal point on Regional Planning (Pianificazione Territoriale). 1996 Graduation (Italian doctor’s degree) with full marks and honours. Title of the Graduation: The Venice Town Council Geographic Information System — Issues of Methodology in the User Requirement Analysis (URA) and of Computer–Assisted Urban Management. The tutor of the thesis was Francesco Gosen, the co–tutor Alberto Giordano. Winner of a one year’s research scholarship within the project “Optimisation of Base Cartography Supporting Items Related to the Venice Lagoon System”. Renunciation of the scholarship. 1997 Inclusion to the list of “Conservators of the Subject” (“cultori della materia”) within the Graduation Course in Urban, Regional, and Environmental Planning at the Venice University of Architecture in the field of Methods and Techniques in Analysis, Representation, and Assessment. 1998 Winner of a scholarship for the collaboration to research activities for the years 1999-2000 within the programme “Knowledge–Building Processes, Urban and Regional Information Systems, Plan–, Project–, and Decision–Making” at the Department of Urban Planning of the Venice University of Architecture. 2000 Co–tutor (with Francesco Sbetti) of the graduation thesis in Urban, Regional, and Environmental Planning entitled “Expecting Places — An Analysis on Disuse Potential in the Province of Treviso, Italy” by Patrizia Del Rosso and Roberto Volpato. Tutor: Domenico Patassini, Venice University of Architecture. 2001 Authorised Dealer for the TNT GIS–products made by MicroImages, Inc., Lincoln, Nebraska, USA. 2002 Member of the National Association of Town, Regional, and Environmental Planners. 2003 Member of the association AM/FM/GIS (Automated Mapping / Facilities Management / Geographic Information Systems) Italia. 2004 Convention with the IUAV University at Venice in order to carry out practical training and vocational orientation courses for university students at the Hedorfer Office. State law examination for the admission to practise the profession of regional planner (“pianificatore territoriale”) passed on the first attempt with 88 of 100 points. Contract lecturer at the IUAV University at Venice for the academic year 2004– 2005.

Languages

German Native language. Italian Perfect active and passive performance of written and

spoken language.

English Good performance, frequently used in work. French Good performance, especially spoken language. Spanish Modest performance. Dutch Reading capability. … Available for learning additional languages.

Information Technology

Operating systems DOS, Microsoft Windows (all versions), UNIX (HP– UX, DEC OS, IBM–AIX, Linux), and X–Windows, Apple–Macintosh.

Programming languages DOS batch–programming, UNIX shell–programming, C/C++, Pascal, Microsoft Windows API, Perl, PHP, Tcl/Tk, Java, Basic, QBasic for Microsoft DOS, Visual Basic, macro languages of TNT (SML), Arc/Info (AML), Arcview (Avenue), AutoCAD (scripting and AutoLisp), Mini-CAD/VectorWorks, MapInfo and products of Microsoft Office. C/C++ API of the TNT Products and Arc/Info. Hypertext Mark–up Language (HTML) and JavaScript. End–user applications Standard office applications (word processing, electronic typesetting, mainly L^{A}TEX, spreadsheet and other DBMSs, publishing, CAD) and geospatial information management software, especially MicroImages (TNT), ESRI (Arc/Info e Arcview), and Autodesk (AutoCAD Map) products. Also SPRING by the Brazilian INPE, CartaLinx by Clark Labs, MapInfo, MicroStation Geographics by Bentley, Intergraph (MGE, I/RAS B and C).

Professional Experiences

Commissions

2000–2001 Numeric encoding of the urban planning instruments and house numbering Comune di (Township of) Duino–Aurisina, Province of Trieste Projecting and implementation of a bilingual (Italian and Slovene) geospatial information system on the management of the municipal planning instruments and house numbers.

2000–2001 Preliminary studies for the Fauna and Hunting Plan Provincia di Venezia (Province of Venice) Preliminary study regarding Regional, Provincial, and Municipal planning instruments, and the state of the territory in general, in order to determine the possibility to establish a set of “protection oases” and “capture and re–population zones” that had yet been identified by the staff of the Protection and Revalorisation Department of the Provincial administration.

1997–1999 GIS for the management of the zoning plan general revision Comune di Venezia (Venice Township) Development of a geospatial information system to allow computer–assisted mangement of the General Programme of Urban Planning (Piano Regolatore Generale). The work has been carried out while two GPUP revision acts regarding the mainland part of the Venice Township (Mestre and Marghera) and the Cavallino Peninsula were prepared.

Consulting

2003 Programming activities within a survey project of paved surfaces Büro für Umwelttechnologie GmbH (BFU, Office for Environmental Technologies Ltd), Gelnhausen / GIS Team, Gießen, Germany Development of software tools to support the electronical encoding activities of built–up and paved surfaces, previously surveyed by employing air photos, and to support the further data processing within the enterprise GIS.

2002 Encoding of contour lines in the Watt Amt für ländliche Räume (ALR, Office for Rural Areas), Husum / GIS Team, Gießen, Germany Quality control and integrative manipulation of a semi–automatic contour line digitisation.

2001–2002 Semi–automatised correction of airphotos Amt für ländliche Räume (ALR, Office for Rural Areas), Husum / GIS Team, Gießen, Germany Within a scanning and georeferencing work: semi–automatised retouching (by employing linear and second–order polynomial interpolation methods) of airphotos in order to remove geographical grid–lines which were yet impressed on the original prints.

1999–2002 Organisation and visualisation of local trade data Comune di (Township of) Loria, Province of Treviso / Venice University of Architecture Development and implementation of a geospatial data base regarding seedling nursery and seedling trade firms in the township district of Bessica, and production of some cartographical compositions with historical contents.

1999–2000 Planning–related spatial analysis and numeric base and thematic cartography Ville de (City of) Bambari, Central African Republic / Venice University of Architecture Development of a geospatial information system in order to manage base data and analytical and planning information at municipal level within the “Municipal development support project — Preliminaries to the urban management plan”.

1998–1999 and 2001–2002 GIS for the management of the zoning plan general revision Comune di (Township of) Castelfranco Veneto, Province of Treviso / Franco Posocco Architect, Venezia–Mestre Participation to the planning team and development of a geospatial data base for the IT–based management of the general regulatory plan during a plan revision operation.

1998 Planning–related spatial analysis and corresponding cartography Comune di (Township of) Villorba, Province of Treviso / Franco Posocco Architect, Venezia–Mestre Structuring, georeferencing, querying, and visualisation of the town council data base on local commercial, industrial, and service activities within the preliminary analysis to the thematic revision of the general regulatory plan concerning industrial settlements.

1996 Environmental numeric cartography and corresponding data base THETIS Consortium Production, within an environmental impact study, of digital base and thematic maps, and organisation of the referring data base.

1994–1996 Participation to the OPEN Project: Information requirements analysis and development of a GIS prototype for Porto Marghera Comune di Venezia (Venice Township) / Consorzio “Venezia Ricerche” (Consortium “Research at Venice”) / Venice University of Architecture Development of a geospatial information system for the Porto Marghera industrial and harbour area as a prototype for a future system involving the whole township territory.

1994–1995 Data input and GIS counselling at the Provincial planning office Provincia di Venezia (Province of Venice) / Consorzio “Venezia Ricerche” (Consortium “Research at Venice”) Insertion of geospatial data concerning the Preliminary Project for the Province–Wide Regional Plan into the Planning Office GIS and consulting tasks on how to develop the spatial planning GIS of the Provincial administration.

Didactics

2004–2005 Contract lecturer within the course of studies in Geographical Information Systems IUAV University at Venice / NETTUNO Consortium Contract lecturer during the academic year 2004–2005 for the course in “Geographical Information Systems 2” (distance university format) within the course of studies in “Geographical Information Systems” totalling 4 CFUs with 30 teaching hours.

2003 Teaching within a vocation–oriented master course IUAV University at Venice Teaching within the “Vocation–Oriented ESF Master in Evaluation of Local Projects and Public Services” for module “Performance Evaluation: Clean–up Scenarios of Contaminated Sites” with a duration of 36 hours. The course was co–financed by the Social Fund of the European Union.

2000 Teaching within a vocational training course ISFID Venezia– Marghera Teaching within the vocational training course as “Environmental Impact Analysis Technician” for module “8b — EIA: EDP–Methodologies” for the duration of 36 hours. The course was co–financed by the Social Fund of the European Union.

1998–1999 Teaching within a vocational training course ENAIP Friuli–Venezia Giulia (Friuli–Venetia Julia) Teaching within the vocational training course as “Urban Renewal Technician” for module “A2.2 — Geographical Information Systems (GIS)” for the duration of 56 hours. The course was co–financed by the Social Fund of the European Union.

1998 University students’ training CEPU Venezia–Mestre Short course of fifteen hours in programming with the Java language at a private organisation specialised in additional university students’ training.

1997–1998 Didactic collaboration Venice University of Architecture Didactic collaboration at the Diploma Course in Geographical Information Systems within the “Laboratory” (1997), by teaching within the courses of Numeric Cartography (1997), Programming Languages (1997), Representation of the Territory and the Environment (1998), and by teaching the students to directly work with GIS software (1998). The commission had been assigned after a competition by qualifications (without exam or interview).

Single Lessons

Contextual knowledge generated by a decision support system for brownfield development: The case of Porto Marghera (Venice, Italy) (with Domenico Patassini) Lesson at the 19th cycle of the PhD programme in Analysis and Governance of Sustainable Development (lesson No. 10 of module 8 “Brownfield Management”), School for Advanced Studies in Venice Foundation, 1st April 2004.

Codifica elettronica di strumenti urbanistici: Castelfranco Veneto e Duino–Aurisina / Devin–Nabrežina Lesson at the course “The Role of GIS Within Town Planning”^{∗} (Prof. Stefano Alonzi) of the Master in Cartography and Geographical Information Systems^{∗}, University of Trieste, 11th November 2003.

Risk and Planning Lesson at the Venice Summer School in Analysis and Governance of Sustainable Development, module “Risk–Based Rehabilitation of Contaminated Megasites”, organised by the Venice International University (VIU), San Servolo Island, Venice 22nd–26th September 2003.

Sistemi informativi geografici: Una panoramica Lesson at the continuation course Spatial Information Management: rules, methods, and instruments (GIS)^{∗} organised by the Interdisciplinary Local Government Research Centre (CISEL), Rimini 22nd–24th April 1998. The lesson has been repeated during the second edition of the continuation course held at Rimini from 15th to 17th December 1998.

University Research

2001–2002 ELGIRA: A knowledge–support procedure for Porto Marghera brownfields at Venice, Italy IUAV University at Venice / Lagoon Research Consortium Development and implementation of a methodological, information, and training model devoted to evaluation of different brownfield development techniques at Porto Marghera, Venice, Italy.

1999–2000 The Venice Metropolitan Area: Infrastructure systems and central places Venice University of Architecture Projecting, implementation, and management of a GIS supporting the research activities on the Venice metropolitan area, with particular attention to interoperability with mobility models and simulations and to new spatial analysis methods.

1996–1997 and 1998–2000 Insertion of planning instruments into different GISs Venice University of Architecture Individuation of methods for computer assisted management of planning instruments, based on GIS technology and with respect to different configurations of the operating platform and base software.

∗All the titles (not the related documents) provided with an asterisk have been translated from their original version into the English language.

1996–1998 Optimisation of base cartography supporting items related to the Venice lagoon system Venice University of Architecture Development and implementation of an organisational model for an experimental geospatial information system within the Venice Lagoon.

Practical Training

1986 Landscape Plan Regionalverband (Region) Neckar–Alb, Tübingen, FR Germany Individuation of regional green belts and green caesurae between settlement agglomerations.

Publications and Contributions

Publications

Contextual Knowledge Generated by a Decision Support System for Brownfield Development: The Case of Porto Marghera (Venice, Italy) (with Domenico Patassini, Paola Cossettini, Enrico De Polignol, Chiara Paneghetti, and Enrico Rinaldi) In course of publication in Beyond Benefit Cost Analysis — Accounting for Non-Market Values in Planning Evaluation, edited by Donald Miller and Domenico Patassini, Ashgate, Aldershot, UK 2004.

Duino–Aurisina / Devin–Nabrežina — Il GIS bilingue per il PRGC / Dvojezični GIS za SORN In course of publication in Archi @ Media.

Castelfranco Veneto — Il GIS per la Costruzione del Nuovo PRG In “Archi @ Media” (2/2001 (April)).

Il Nuovo Piano Regolatore per Mestre — Struttura del Piano e Codifica Numerica In “Archi @ Media” (1/2000 (March)).

Informazioni Urbanistiche nei GIS In “Procedure Digitali per la Pianificazione Ambientale”, edited by Igor Jogan and Domenico Patassini, 165–201, Il Rostro, Milan, Italy 2000.

Usefulness of GIS in Strategic Planning In “Marittima Workshop — G.I.S. and Strategic Planning”, edited by Giuseppe Longhi, 53–70, Venice University of Architecture, Department of Urban Planning, October 1999.

Il GIS per la Gestione degli Strumenti Urbanistici del Comune di Venezia (with Francesco Contò, Massimo Mazzanti, and Paolo Barbieri) In “Geomedia” (1/1998 (February)). The same article, with few changes, has also been published on the conference proceedings of the itinerant conference GIS Itinera 1997 and is referred to the contributions of Hedorfer, Contò, and Mazzanti held at the Venice (31st October 1997) and Perugia (18th November 1997) editions.

Le monografie — 1. Speciale GIS (with Pier Francesco Ricci) supplement to the review “l’ufficio tecnico”, (3, March 1997). Chapters: Geographical Information Systems: An Overview For Everybody^{∗}; GIS Software^{∗}; Urban and Regional Planning^{∗}. Reprint in the numbers 1–6/1998 and 1/1999 of the journal “Geomedia” as a column entitled The GIS Tutorial^{∗}.

Requisiti informativi e gestione informatica degli strumenti urbanistici: il SIT del Comune di Venezia In “Cronache Ca’ Tron” (internal journal of the Venice University of Architecture) (8/1996). Summary of the Graduation Thesis.

Conference Participations

Un modèle structurel pour métadonnées (with Alberta Bianchin) Contribution to the 3^{èmes} Journées Cassini, Marne–la–Vallée, Seine–et–Marne, France, 26th–27th November 1998.

Strutturazione e Trattamento dei Metadati nel SISALV (with Alberta Bianchin) In “Rilevamento, rappresentazione e gestione dei dati territoriali e ambientali — Atti della 2^{a} Conferenza Nazionale delle Associazioni Scientifiche per le Informazioni Territoriali e Ambientali (ASITA) 24–27 novembre 1998” (conference proceedings), 2 voll., 113–117 (volume 2), ASITA, Bolzano, Italy 1998.

Il GIS per la Gestione degli Strumenti Urbanistici del Comune di Venezia (with Francesco Contò, Massimo Mazzanti, and Paolo Barbieri) Contribution to the itinerant conference GIS Itinera 1997 at the Venice (31st October 1997) and Perugia (18th November 1997) editions.

Publications and Contributions within the ELGIRA Project (since 2002)

Valutazione di progetti di bonifica dei siti inquinati. El.GI.R.A. — Una procedura di aiuto alla conoscenza nelle aree di bonifica di Porto Marghera (Venezia) (with Paola Cossettini, Enrico De Polignol, Chiara Paneghetti, Domenico Patassini, and Enrico Rinaldi) Contribution to input 2003 — terza conferenza nazionale su informatica e pianificazione urbana e territoriale: «costruzione e gestione della conoscenza», Pisa, Italy 5th–7th June 2003.

ELGIRA — Modello valutativo per la gestione degli scenari di bonifica a Porto Marghera (Venezia) (with Paola Cossettini, Enrico De Polignol, Chiara Paneghetti, Domenico Patassini, and Enrico Rinaldi) Contribution to the VI Conferenza Nazionale dell’Associazione Italiana di Valutazione (AIV), Reggio di Calabria, Italy 10th–11th April 2003.

Hints on context knowledge within a support system for brownfield development: the case of Porto Marghera (Venice) (with Paola Cossettini, Enrico De Polignol, Chiara Paneghetti, Domenico Patassini, and Enrico Rinaldi) Contribution to the Fifth International Workshop on “Evaluation in Planning”, Venice, Italy 14th–15th February 2003.

ELGIRA — A knowledge support procedure for rehabilitating contaminated real estate in Porto Marghera (Venice) (with Enrico De Polignol, Chiara Paneghetti, Domenico Patassini, and Enrico Rinaldi) Contribution to the 1st Iuav Conference on: Brownfields as Opportunities for Sustainable Development, Venice, Italy 31th January – 1st February 2003.

Valutazione di progetti di bonifica dei siti inquinati — Una procedura di aiuto alla conoscenza di contesto: El.GI.R.A. (with Chiara Paneghetti, Domenico Patassini, and Enrico Rinaldi) Contribution to the V Conferenza Nazionale dell’Associazione Italiana di Valutazione (AIV), Bologna, Italy 9th–11th May 2002.

Publications and Contributions within the SISALV Project (1997– 1999, see at page 7)

Utilisation intégrée de données cartographiques et de simulations de phénomènes de pollution: une proposition pour la Lagune de Venise (with Alberta Bianchin, and Donatella Schiuma) Poster presented at the conference ICA 1999, Ottawa, Canada, 16th–20th August 1999.

The Venice Lagoon Experimental GIS at the IUAV (with Alberta Bianchin, and Donatella Schiuma) Poster presented at the 2^{nd} AGILE (Association of Geographic Information Laboratories in Europe) Conference on Geographic Information Science, Rome, Italy, 15th–17th April 1999.

Il GIS Sperimentale presso l’IUAV per la Laguna Veneta (with Alberta Bianchin, and Donatella Schiuma) In “Rilevamento, rappresentazione e gestione dei dati territoriali e ambientali — Atti della 2^{a} Conferenza Nazionale delle Associazioni Scientifiche per le Informazioni Territoriali e Ambientali (ASITA) 24–27 novembre 1998” (conference proceedings), 2 voll., 329–334 (volume 1), ASITA, Bolzano, Italy 1998.

Integrazione di un modello di qualità dell’acqua in ambiente GIS: applicazione alla Laguna di Venezia (with Alberta Bianchin, Roberto Pastres, Cosimo Solidoro, and Vittorio E. Brando) Poster presented at the 2^{a} Conferenza Nazionale delle Associazioni Scientifiche per le Informazioni Territoriali e Ambientali (ASITA), Bolzano, Italy 24th–27th November 1998.

SISALV — Venice Lagoon Experimental GIS at the IUAV — GIS Sperimentale presso l’IUAV per la Laguna Veneta (with Alberta Bianchin, Roberto Pastres, Vittorio E. Brando, Enrico Rinaldi, Donatella Schiuma, Cosimo Solidoro, and Sandro Soramaè) Poster presented at the International Conference on Education and Training in Integrated Coastal Area Management: The Mediterranean Prospect, Genoa, Italy, 25th–29th May 1998 and at the 2^{a} Conferenza Nazionale delle Associazioni Scientifiche per le Informazioni Territoriali e Ambientali (ASITA), Bolzano, Italy 24th–27th November 1998.

SISALV 1997 — Sistema Informativo Geografico Sperimentale presso l’IUAV per la Laguna Veneta Contribution to the itinerant conference GIS Itinera 1997 at the Venice (31st October 1997) edition.

Minor Contributions

Beyond Benefit Cost Analysis — Accounting for Non-Market Values in Planning Evaluation edited by Donald Miller and Domenico Patassini. Ashgate, Aldershot, UK 2004 (in course of publication). Technical editing of the book.

Villici industrianti commercianti — Le radici storiche e culturali di una vicenda di sviluppo locale. Il caso di Bessica, Villa dell’Alto Trevigiano by Augusto Cusinato. Franco Angeli, Milan, Italy 2003. See “Organisation and visualisation of local trade data” at page 4.

Software Development and Documentation

Italian version of the TNT Products Italian linguistic adaptation (translation, localisation, and maintenance) of the GIS software product familiy “TNT Products”, produced by MicroImages, Inc., Lincoln, Nebraska, USA. Available since version 6.30 released in April 2000..Further information is available on the Internet at address http://www.microimages.com/i18n/locales/italian/ v680.htm.

Prodotti TNT — L’analisi geospaziale diventa facile Italian translation of the English language publication TNT Products — Geospatial Analysis Made Easy, edited and published by MicroImages, Inc., Lincoln, Nebraska, USA 2000 (2002), 28 pages. The document is available on the Internet at address ftp://ftp.microimages.com/pub/outgoing/fordealers/_it_italian/bro chure-ITA.pdf .

Cartografia numerica: Creazione di carte topografiche Italian translation of the English language publication Digital Cartography: Making Topographic Maps by Randall B. Smith, MicroImages, Inc., Lincoln, Nebraska, USA 2002 (2002), 12 pages. The document is available on the Internet at address ftp://ftp.microimages.com/pub/outgoing/fordealers/_it_italian/ita_to pomap.pdf .

Other Information

Public Institutions

Member of the Urban Planning (1991–1997) and of the Public Works (1995– 1997) Commission of the Venice Township District n. 11. Member of the Building Commission of the Township of Mirano (Province of Venice, 1995–1996). Member of the Managing Board of the ATER (Azienda Territoriale per l’Edilizia Residenziale, Area Enterprise for [Public] Housing) of the Province of Venice as representative of the Venice Township (1997–1999).

Actual Contracts

My actual contract obligations concern only the research work at the IUAV on brownfield development issues at Porto Marghera and the teaching activity at the course of studies in GIS.